OpenStudy (anonymous):

solve 2x-8/4=5 7/2 -7/2 14 -14

OpenStudy (tommo_lcfc):

Is it \[2x - 8/4 = 2x - 2\] or \[(2x-8)/4\]? The former produces 2x - 2 = 5 so x = 14. The latter produces 2x - 8 = 20 (multiplying up) so x = 14 again.